# Survey on Modeling of Articulated Objects

### v1

@@ 6 Future Discussion and Conclusion
Generalization across objects with varying structures. Handling the diversity of articulated objects with large variations in the geometric and kinematic structures is one of the major challenges in the field. Existing methods are often designed for specific object categories or assume a certain level of prior knowledge about the object is available to the system. Given limited training data and the high cost of data collection, it is important to develop methods that can generalize to novel objects in a data-efficient manner.
Improved accuracy and robustness in the perception models. Although significant progress has been made in the development of articulated part perception models, there is still a need for more accurate and robust methods that can handle occlusions, clutter, and noise in the input data. This is particularly important for applications such as robotics and autonomous driving, where the models need to be able to handle real-world scenarios.
Incorporating physical constraints into the modeling process. Articulated objects are subject to physical constraints that are often not considered in the existing modeling process. Examining the physical constraints of the objects can help to improve the accuracy of the articulation modeling and enable the generation of physically plausible animations and simulations. This consideration is also useful in articulated object creation for applications such as virtual reality, where the goal is to create realistic and physically plausible models.
@@
Detailed interior and part geometry modeling. Modeling interior and part-level geometry in detail for articulated objects is an important aspect of geometric modeling that has not been fully explored. For the object reconstruction task, the quality of the models is often limited by the quality of the input observation which is often partial due to view occlusion and sensor noises. Developing methods that can handle these challenges and generate detailed and accurate 3D models is an important direction for future research.
Generative AI for articulated objects. The recent advances in generative models and large language models have shown great potential in various computer vision and graphics tasks. Applying these techniques to the domain of articulated object modeling can be a promising direction to provide a better understanding and accessibility of the complex 3D articulated structures.
@@
Conclusion. This survey systematically reviewed the progress in 3D modeling for articulated objects, with a focus on articulated part perception and articulated object creation. By examining the development of articulated object modeling across the two axes of geometric processing and articulation modeling, we have identified significant advancements and ongoing challenges in the field. We also highlighted the potential research directions that can drive future progress in the field. We hope that this survey serves as a foundational reference for researchers and practitioners in computer vision and graphics, offering insights into the complexities of articulated object modeling and inspiring new research in this area.
@@
Acknowledgements. This work was funded in part by a Canada Research Chair and an NSERC Discovery grant. We thank Angel X. Chang and Hao (Richard) Zhang for helpful discussions and feedback on preliminary drafts.

### v3

@@ 6 Discussion and Conclusion
Generalization across objects with varying structures. Handling diverse articulated objects with large variations in geometric and kinematic structures is one of the major challenges in the field. Existing methods are often designed for specific object categories or assume a certain level of prior knowledge about the object is available to the system (e.g., the number of parts, which joint types are considered). Given limited training data and the high cost of data collection, it is important to develop methods that can generalize to novel objects in a data-efficient manner. A promising direction is to explore the use of few-shot learning, meta-learning, and domain adaptation techniques to enhance the versatility of articulated object understanding. Recent advances in large language models, vision-language models, and multimodal large language models can also be powerful prior to integrate into articulated object perception and manipulation pipelines, driving improvements in the generalization and robustness of the models.
Improved robustness in the perception models. While significant progress has been made in articulated part perception, more accurate and robust methods are needed to handle occlusions, clutter, and noise in input data. Most existing approaches focus on object-level perception with clear, object-centric observations, whereas perceiving articulated objects in complex scenes with multiple objects and partial observations remains an open challenge delitzas2024scenefun3d. Addressing this gap is essential for making these methods practical and deployable in real-world applications, including functionality understanding, affordance learning, and robotic manipulation, where models must reliably interpret and interact with articulated objects in unstructured environments.
@@
Detailed interior and part geometry modeling. Modeling interior and part-level geometry in detail for articulated objects and further composing such detailed interactable objects at the scene level is an important aspect of geometric modeling that has not been fully explored. For the object reconstruction task, the quality of the models is often limited by the quality of the input observation which is often partial due to view occlusion and sensor noise. Developing methods that can handle these challenges and generate intricate and accurate 3D models of the articulated object (e.g., fully functional cars, fridges with shelves and containers allowing small objects to be placed inside) is another promising direction for future research.
Generative AI for articulated objects. The recent advances in articulated object generation have shown great potential in creating realistic and diverse 3D models of articulated objects. Due to the highly structured nature and complexity in both geometry and articulation at the part level, generative models for articulated objects face unique challenges compared to general object generation. The existing methods either prioritize structure modeling or geometry modeling, and the integration of both aspects while maintaining kinematic plausibility remains an open challenge. Key challenges include designing efficient representations with compatible generative paradigms to capture the multi-faceted nature of articulated objects, and developing controllable models that can be prompted with different user inputs. Additionally, data scarcity limits progress, as high-quality datasets with detailed part geometry and articulation annotations are rare, making it difficult to train scalable and generalizable models. Promising future directions include expanding datasets, improving the representation learning, enhancing the multi-modal generation capabilities, and advancing controllable generative models for interactive design. Another interesting venue is exploring domain-specific language models for articulated objects in combination with large language models. The compositionality and the hierarchical structure of the articulated objects make them well suited for code- or program-based descriptions, and the integration of language priors can potentially improve the generation diversity and the model interpretability. Ultimately, integrating generative AI with simulation environments will be essential for advancing interactive applications and embodied AI research, enabling more realistic modeling and interaction with everyday articulated objects.
Incorporating physical constraints into the modeling process. Articulated objects are subject to physical constraints (e.g., collisions, materials) and physical properties (e.g., mass, inertia), which are often overlooked in existing modeling approaches. Most current generative and perception models focus on capturing articulation geometry and motion patterns but do not explicitly enforce real-world physics, leading to unrealistic or non-functional outputs. Integrating the physical constraints of the objects can help to improve the accuracy of the articulation modeling and ensure that generated objects not only look realistic but also move and interact in a physically plausible manner. Considering physical constraints is particularly crucial for simulation-driven applications such as robotic manipulation, digital twins, and physics-based animation, where real-world interaction dynamics must be accurately replicated. This consideration is also useful in articulated object creation for applications such as virtual reality and gaming, where objects must react naturally to user interactions, maintaining realistic movements, and collision handling. Future research should explore integrating physics engines directly into the modeling pipeline to enforce kinematic and dynamic feasibility during object creation. Additionally, learning-based approaches could benefit from differentiable physics simulation which allows models to be trained with explicit physics constraints, improving both realism and generalization. By embedding physical reasoning into the content creation process, we can bridge the gap between visual realism and functional accuracy, paving the way for more deployable and interactive articulated object models in robotics and virtual environments.
Leveraging human-in-the-loop data. Everyday human interactions with articulated objects produce a large amount of data in the form of long natural videos, egocentric videos, and human demonstrations (see examples in Figure 17). These data are rich in articulated object information but also present challenges such as fast-moving objects, limited field-of-view camera viewpoints, streaming setups where the parts may disappear and reappear, and lots of occlusions and dynamics due to humans/hands in the view. Collecting and using such data for articulated object modeling can enable generalization to diverse real objects and interactions.
Evaluation and benchmarking. Due to the multi-faceted nature of articulated objects, the evaluation of the quality of the created objects is currently not standardized. The challenges of evaluation are partially due to the lack of ground truth data and the complexity of the task. And even when the ground truth annotation is available, the metrics that have been proposed by prior work have flaws and are not fully and clearly specified. There is a need for more systematic evaluation metrics that can evaluate different aspects of articulated object creation, including geometry, articulation, and part relationships, and also an object as a whole in the context of a scene where interaction can take place.
@@
Conclusion. This survey systematically reviewed the progress in 3D modeling for articulated objects, with a focus on articulated part perception and articulated object creation. By examining the development of articulated object modeling across the two axes of geometric and articulation modeling, we have identified significant advances and ongoing challenges in the field. We also highlighted potential research directions that can drive future progress in the field. We hope that this survey serves as a foundational reference for researchers and practitioners in computer vision and graphics, offering insights into the complexities of articulated object modeling and inspiring new research in this area.
@@
Acknowledgments. This work was funded in part by NSERC Discovery grants (RGPIN-06489-2019, RGPIN-2022-03111), a Canada Research Chair grant (CRC-2019-00298), and a startup grant from Simon Fraser University (N000449). We thank Angel X. Chang, Richard Zhang, and Han-hung Lee for their helpful comments and discussions.
[…]